OpenStudy (anonymous):

Find the zero : y=36x^2+36x+9 Ugh how to solve this one. :"(

OpenStudy (zpupster):

try and reduce it first div by 9

OpenStudy (anonymous):

so 4x+4+1?

OpenStudy (zpupster):

not quite 4x^2 + 4x +1 now factor

OpenStudy (anonymous):

ah i don't know how to factor that

OpenStudy (anonymous):

What is the common factor?

OpenStudy (anonymous):

And highest.

OpenStudy (anonymous):

2 and 4?

OpenStudy (zpupster):

ok I will show you one wayS: (2x )(2x ) first term mult = 4x^2 (2x 1 )(2x 1 ) our last term is +1 it (-1)(-1) or(1) (1) we know our middle term is 4x=2x+2x we get this by mult outside terms and inside terms so (2x+1)(2x+1) is our factor

OpenStudy (anonymous):

is there any shortcut method

OpenStudy (anonymous):

i can't understand that one. how'd u got 2x?

OpenStudy (zpupster):

well if you have x^2 you can see it is x times x 4x^2 = 2x

OpenStudy (anonymous):

how about the remaining 4x+1 ?

OpenStudy (zpupster):

I am using the Foil method in reverse multiply the first outside inside last so I have to figure out the first terms 2x times 2x is a good bet because it equals 4x^2

OpenStudy (anonymous):

oh okaaay thanks a lot

OpenStudy (zpupster):

did you get how to find the zero

OpenStudy (anonymous):

-1/2 ?

OpenStudy (anonymous):

is that correct?

OpenStudy (lncognlto):

Yes, if 2x+1 = 0, then x = -1/2

OpenStudy (lncognlto):

@MCath

OpenStudy (zpupster):

good job